# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v2

@@ 4.2 Finitely many narrow wedges and approximations
Given $\mathfrak{h}\in m\mathrm{NW}_{0}$. Let $(\mathfrak{h}^{\varepsilon})_{\varepsilon>0}$ be the approximating sequence of initial height functions for TASEP defined as in (4.9) and (4.10). Given $z_{1},\cdots,z_{m}\in\mathbb{C}$ with $|z_{i}|=r<1$ for $1\leq i\leq m-1$. To lighten the notation we will suppress the dependency on $\varepsilon$ at most places and write
$\mathcal{D}_{Y^{\varepsilon}}(z_{1},\cdots,z_{m-1}):=\mathcal{D}_{Y(\mathsf{h}^{\varepsilon})}(z_{1},\cdots,z_{m-1};\mathbf{k}^{\varepsilon},\mathbf{a}^{\varepsilon},\mathbf{t}^{\varepsilon}),$
where $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})=\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1};\mathbf{k},\mathbf{a},\mathbf{t})$ is defined in Section 3.2.1. Here we use boldface letters to denote vectors, for example, $\mathbf{k}:=(k_{1},\cdots,k_{m})$. Assume the parameters satisfy
@@
$\lim_{\varepsilon\to 0}\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathcal{D}_{Y^{\varepsilon}}(z_{1},\cdots,z_{m-1})=\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1}).$
@@
Let $\mathcal{D}^{(\mathbf{n})}_{Y^{\varepsilon}}$ and $\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}$ be as in (3.17) and (2.21), where $\mathfrak{h}$ is given by (4.8) and $Y^{\varepsilon}=Y(\mathsf{h}^{\varepsilon})$ is described in (4.9). Then for each $\mathbf{n}\in(\mathbb{Z}_{\geq 0})^{m}$ and $(z_{1},\cdots,z_{m-1})\in(\mathbb{D}(0,1))^{m-1}$, we have
$\lim_{\varepsilon\to 0}\mathcal{D}^{(\mathbf{n})}_{Y^{\varepsilon}}(z_{1},\cdots,z_{m-1};\mathbf{k}^{\varepsilon},\mathbf{a}^{\varepsilon},\mathbf{t}^{\varepsilon})=\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1};\bm{\alpha},\bm{\beta},\bm{\tau}).$ (4.12)
@@
$\left|\mathcal{D}^{(\mathbf{n})}_{Y^{\varepsilon}}(z_{1},\cdots,z_{m-1};\mathbf{k}^{\varepsilon},\mathbf{a}^{\varepsilon},\mathbf{t}^{\varepsilon})\right|\leq\prod_{\ell=1}^{m-1}\frac{(1+|z_{\ell+1}|)^{2n_{\ell+1}}}{|z_{\ell}|^{n_{\ell+1}}|1-z_{\ell}|^{n_{\ell+1}-n_{\ell}}}\cdot\prod_{\ell=1}^{m}n_{\ell}^{n_{\ell}}\cdot C^{n_{1}+\cdots+n_{m}},$ (4.13)
for any $\mathbf{n}=(n_{1},\cdots,n_{m})\in(\mathbb{Z}_{\geq 0})^{m}$ and $(z_{1},\cdots,z_{m-1})\in(\mathbb{D}(0,1))^{m-1}$.
@@ 4.2.1 Pointwise convergence of the characteristic function
For the approximating sequence of height functions $\mathsf{h}^{\varepsilon}$ described in (4.10), we denote $Y^{\varepsilon}$ the corresponding particle configurations for TASEP. It consists of exactly $M$ clusters of densely packed particles. To lighten the notation we denote temporarily the indices of the rightmost particle of each cluster by $\mathsf{t}_{0},\cdots,\mathsf{t}_{M-1}$, from right to left. We have
@@
$\mathbb{P}(\tau\notin\{\mathsf{t}_{0},\cdots,\mathsf{t}_{M-1}\})=0.$
Here $\tau$ is defined as in (3.4) and the indices $\mathsf{t}_{0},\cdots,\mathsf{t}_{M-1}$ are as in (4.14). Hence
@@
where we are using the fact that $\mathsf{t}_{k}\geq 0$ and $y_{\mathsf{t}_{0}}\geq y_{\mathsf{t}_{k}}-\delta+1$ for all $y_{\mathsf{t}_{k}}-y_{\mathsf{t}_{0}}+1\leq\delta$. Thus we conclude that the right-hand side of (4.23) is bounded above by
@@
Using the bound from (4.28) in (4.26) and summing over $k$, we arrive at the desired estimate (4.18).
@@
(a) Now we prove part (a). We start with rewriting $\chi_{\mathfrak{h}}(\eta,\xi)$ under the assumption that $\mathfrak{h}\in m\mathrm{NW}_{0}$, recall the definition of $\chi_{\mathfrak{h}}(\eta,\xi)$ from (2.9). For $\mathrm{supp}(\mathfrak{h})=\{\omega_{0},\cdots,\omega_{M-1}\}$ with $0=\omega_{0}>\cdots>\omega_{M-1}$, one has $\mathbb{P}(\bm{\tau}_{+}\neq 0)=0$ where $\bm{\tau}_{+}$ is defined in (2.16). Thus it is easy to check that the first and third term on the right-hand side of (2.9) are both equal to $\frac{\mathrm{e}^{\mathfrak{h}(0)(\eta-\xi)}}{\eta-\xi}$ and they cancel each other. On the other hand since $\mathbb{P}(\bm{\tau}_{-}\notin\{\omega_{0},\cdots,\omega_{M-1}\})=0$, the second term on the right-hand side of (2.9) is given by
@@
In order to rigorously show the above convergence, we need to show that (4.38) is uniformly bounded and the dominated convergence theorem applies. Note that the right hand side of (4.38) is the same as that of (4.21). Therefore (4.19) gives a uniform bound for (4.38). This completes the proof. ∎
@@ 4.3 Convergence of the series expansion
In this section we prove Lemma 4.3 and 4.4. We will make the additional assumption that $0<\tau_{1}<\cdots<\tau_{m}$ to make the presentation lighter. The convergence results and arguments in this section still work if $\tau_{i}=\tau_{i+1}$ for some $i$ but the contours need to be chosen carefully to make sure that the integrand has the desired super-exponential decay. Alternatively one can directly work with the limiting KPZ fixed point formula which is continuous with respect to the limit $\tau_{i+1}\to\tau_{i}$ and our choices of the angles in Figure 1 guarantees the convergence of the formula (2.21) even when some time parameters are equal.
@@
Thus for such $\eta$ we have
@@
For the first part of the claim, note that the integrand on the right-hand side of (4.50) converges pointwise to the integrand on the right-hand side of (2.21) by Lemma 4.6(a) and Proposition 4.2(a). On the other hand the cubic exponential decay bound for $|\mathrm{f}_{i}^{\varepsilon}|$ from Lemma 4.6(b) and the quadratic exponential growth estimate (4.19) implies that the integrand on the right-hand side of (4.50) has a cubic exponential decay in every variable $\eta_{i_{\ell}}^{(\ell)},\xi_{i_{\ell}}^{(\ell)}$ as they go to $\infty$, uniform in $\varepsilon$. Thus the dominated convergence theorem applies and the first claim is proved.
For the second part of the claim, note that on $A^{c}$, at least one of the variables, say $\xi_{1}^{(1)}$, lies outside of $\widehat{\Omega}_{0}^{\varepsilon}$. Then Lemma 4.6(b) implies that $|\mathrm{f}_{1}^{\varepsilon}(\xi_{1}^{(1)})|\leq C\exp(-c\varepsilon^{-\frac{3}{2}})$ for all $\xi_{1}^{(1)}\in{\Gamma}_{1,\mathrm{L}}^{\varepsilon}\backslash\widehat{\Omega}_{0}^{\varepsilon}$. On the other hand Proposition 4.2(b) implies that
@@
for some $c,C>0$ and all $1\leq j\leq n_{1}$. The other parts of the integrand remain bounded. Thus
@@
This proves the second part of the claim and Lemma 4.3 follows.
@@
for some constant $c>0$ independent of $\varepsilon$ and $\bm{\xi}^{(1)},\bm{\eta}^{(1)}$. Thus by Hadamard’s inequality
@@
for $1\leq\ell\leq m-1$. Here $D$ is chosen to be the reciprocal of the minimal distance between different $\Gamma^{\varepsilon}$ contours, which can be made postive. Thus
@@ 5 From multiple narrow wedges to compactly supported initial conditions
In this section we extend the multipoint formula (1.6) to any $\mathfrak{h}\in\mathrm{UC}_{c}$ using a density argument. We choose to work at the level of the KPZ fixed point formula which enjoys more symmetry and nicer decay properties.
@@ 5.1 Density and Approximation
Our starting point is the following proposition asserting that our formula is continuous with respect to the initial condition on the space $\mathrm{UC}_{c}$.
@@
for all $\xi\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$. Recall the kernel $\chi_{\mathfrak{h}}$ was defined in Definition 2.2, equation (2.9). Consequently
$\lim_{n\to\infty}\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}^{n}}(z_{1},\cdots,z_{m-1})=\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1}).$ (5.2)
@@
This completes the proof of (5.1). (5.2) follows from the dominated convergence theorem and a similar bound as in Lemma 4.4 whose proof is almost identical and we omit it here. ∎
To extend (1.6) to all $\mathfrak{h}\in\mathrm{UC}_{c}$ using Propsition 5.1, we also need the following proposition asserting that the space $m\mathrm{NW}$ is dense in $\mathrm{UC}_{c}$.
@@
It suffices to consider the restrictions of the functions on [L,L][-L,L]. For each $n$ consider the dyadic intervals $I_{n,k}:=[\frac{k}{2^{n}}\cdot L,\frac{k+1}{2^{n}}\cdot L]$, for $k=-2^{n},-2^{n}+1,\cdots,2^{n}-1$. On each interval $I_{n,k}$ the maximum of $\mathfrak{h}$ exists, let $m_{n,k}\in I_{n,k}$ be one of the argmax, and set
@@
Therefore $\mathfrak{h}^{n}\to\mathfrak{h}$ locally in $\mathrm{UC}$, and hence globally since $\mathfrak{h}^{n},\mathfrak{h}$ are supported in [L,L][-L,L]. ∎
@@
Recall that what we have shown in Proposition 5.2 is that the formula (1.6) holds for initial condition $\mathfrak{h}\in m\mathrm{NW}_{0}$. We would like to prove it for $\mathfrak{h}\in m\mathrm{NW}$ by a shift argument. To this end it is more convenient to have a more general version of the characteristic function $\chi_{\mathfrak{h}}$, defined through (2.15), instead of the original (2.9). The goal of this section is to prove the equivalence of the two, i.e., Proposition 2.4. To this end, we denote
@@
$+\mathrm{e}^{\omega\eta^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{+s\eta}\cdot\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(-\bm{\tau}^{\omega}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}^{\omega}_{+})\xi\right)\mathbf{1}_{\bm{\tau}^{\omega}_{+}<\infty}\right]$
$-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(\bm{\tau}_{-}^{\omega}\eta^{2}+\mathbf{B}(\bm{\tau}^{\omega}_{-})\eta-\bm{\tau}_{+}^{\omega}\xi^{2}-\mathbf{B}(\bm{\tau}^{\omega}_{+})\xi\right)\mathbf{1}_{|\bm{\tau}^{\omega}_{\pm}|<\infty}\right]$
@@
where $\omega_{1}>\cdots>\omega_{k}$. To lighten the notation we will fix $\eta,\xi$ and denote
@@
$\chi_{\mathfrak{h}}^{\omega,1}(\eta,\xi)=\mathrm{e}^{-\omega\xi^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\int_{\mathbb{R}}\mathop{}\!\mathrm{d}b_{1}\,\mathrm{e}^{-s\xi}\mathrm{p}_{\omega-\omega_{1}}(b_{1}-s)\,E^{-}(\omega_{1},b_{1})=\mathrm{e}^{-\omega_{1}\xi^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}b_{1}\,\mathrm{e}^{-b_{1}\xi}\,E^{-}(\omega_{1},b_{1})=\chi_{\mathfrak{h}}^{\omega_{1},1}(\eta,\xi),$
@@
Recall that $\mathrm{p}_{t-s}(x-y)$ is the transition density of a Brownian motion with diffusivity constant $2$. On the other hand
@@
$+\mathbf{1}_{\omega=\omega_{i+1}}\frac{\mathrm{e}^{\omega_{i+1}(\eta^{2}-\xi^{2})+\theta_{i+1}(\eta-\xi)}}{\eta-\xi}.$
Combine (5.5), (5.6) and (5.7) we conclude that
@@
for all $\omega\in[\alpha_{i+1},\alpha_{i})$. Thus by induction $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}$ for all $\omega\in\mathbb{R}$ and $\mathfrak{h}\in m\mathrm{NW}$.
For general $\mathfrak{h}\in\mathrm{UC}_{c}$ supported on [L,L][-L,L], bounded above by $\beta$, we use Proposition (5.2) to find a sequence $\{\mathfrak{h}^{n}\}_{\geq 1}\subset m\mathrm{NW}$ bounded above by $\beta$ whose support are contained in [L,L][-L,L] , such that $\mathfrak{h}^{n}\to\mathfrak{h}$ in $\mathrm{UC}$. Then by Proposition 5.1 we know $\chi_{\mathfrak{h}^{n}}\to\chi_{\mathfrak{h}}$ as $n\to\infty$. A minor variant of Proposition 5.1 with the same proof shows that $\chi^{\omega}_{\mathfrak{h}^{n}}\to\chi^{\omega}_{\mathfrak{h}}$ as $n\to\infty$ for any $\omega\in\mathbb{R}$ as well. Thus
@@
Recall that in Proposition 4.2, we have shown Theorem 1.2 for $\mathfrak{h}\in m\mathrm{NW}_{0}$. We first extend it to $\mathfrak{h}\in m\mathrm{NW}$ using Propsition 2.4. There exists $\alpha,\beta\in\mathbb{R}$ such that $\mathfrak{h}^{\alpha,\beta}:=\mathfrak{h}(\cdot+\alpha)+\beta\in m\mathrm{NW}_{0}$. Thus by the invariance property of the KPZ fixed point (see, e.g., [MQR21][Theorem 4.5]) we have
@@
Apply Proposition 4.2 to the right-hand side of (5.8) and compare the resulting formula with the right-hand side of (1.6) , we see that in order for Theorem 1.2 to hold for $\mathfrak{h}\in m\mathrm{NW}$, it suffices to show that

### v3

@@ 4.2 Finitely many narrow wedges and approximations
Given $\mathfrak{h}\in m\mathrm{NW}_{0}$. Let $(\mathfrak{h}^{\varepsilon})_{\varepsilon>0}$ be the approximating sequence of initial height functions for TASEP defined as in (4.9) and (4.10). Given $z_{1},\ldots,z_{m}\in\mathbb{C}$ with $|z_{i}|=r<1$ for $1\leq i\leq m-1$. To lighten the notation we will suppress the dependency on $\varepsilon$ at most places and write
$\mathcal{D}_{Y^{\varepsilon}}(z_{1},\ldots,z_{m-1}):=\mathcal{D}_{Y(\mathsf{h}^{\varepsilon})}(z_{1},\ldots,z_{m-1};\mathbf{k}^{\varepsilon},\mathbf{a}^{\varepsilon},\mathbf{t}^{\varepsilon}),$
where $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})=\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1};\mathbf{k},\mathbf{a},\mathbf{t})$ is defined in Section 3.2.1. Here we use boldface letters to denote vectors, for example, $\mathbf{k}:=(k_{1},\ldots,k_{m})$. Assume the parameters satisfy
@@
$\lim_{\varepsilon\to 0}\prod_{\ell=1}^{m-1}\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathcal{D}_{Y^{\varepsilon}}(z_{1},\ldots,z_{m-1})=\prod_{\ell=1}^{m-1}\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}).$
@@
Let $\mathcal{D}^{(\mathbf{n})}_{Y^{\varepsilon}}$ and $\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}$ be as in (3.18) and (2.21), where $\mathfrak{h}$ is given by (4.8) and $Y^{\varepsilon}=Y(\mathsf{h}^{\varepsilon})$ is described in (4.9). Then for each $\mathbf{n}\in(\mathbb{Z}_{\geq 0})^{m}$ and $(z_{1},\ldots,z_{m-1})\in(\mathbb{D}(0,1))^{m-1}$, we have
$\lim_{\varepsilon\to 0}\mathcal{D}^{(\mathbf{n})}_{Y^{\varepsilon}}(z_{1},\ldots,z_{m-1};\mathbf{k}^{\varepsilon},\mathbf{a}^{\varepsilon},\mathbf{t}^{\varepsilon})=\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1};\bm{\alpha},\bm{\beta},\bm{\tau}).$ (4.12)
@@
$\left|\mathcal{D}^{(\mathbf{n})}_{Y^{\varepsilon}}(z_{1},\ldots,z_{m-1};\mathbf{k}^{\varepsilon},\mathbf{a}^{\varepsilon},\mathbf{t}^{\varepsilon})\right|\leq\prod_{\ell=1}^{m-1}\frac{(1+|z_{\ell+1}|)^{2n_{\ell+1}}}{|z_{\ell}|^{n_{\ell+1}}|1-z_{\ell}|^{n_{\ell+1}-n_{\ell}}}\cdot\prod_{\ell=1}^{m}n_{\ell}^{n_{\ell}}\cdot C^{n_{1}+\cdots+n_{m}},$ (4.13)
for any $\mathbf{n}=(n_{1},\ldots,n_{m})\in(\mathbb{Z}_{\geq 0})^{m}$ and $(z_{1},\ldots,z_{m-1})\in(\mathbb{D}(0,1))^{m-1}$.
@@ 4.2.1 Pointwise convergence of the characteristic function
For the approximating sequence of height functions $\mathsf{h}^{\varepsilon}$ described in (4.10), we denote $Y^{\varepsilon}$ the corresponding particle configurations for TASEP. It consists of exactly $M$ clusters of densely packed particles. To lighten the notations, we denote temporarily the indices of the rightmost particle of each cluster by $\mathsf{t}_{0},\ldots,\mathsf{t}_{M-1}$, from right to left. We have
@@
$\mathbb{P}(\tau\notin\{\mathsf{t}_{0},\ldots,\mathsf{t}_{M-1}\})=0.$
Here $\tau$ is defined as in (3.4) and the indices $\mathsf{t}_{0},\ldots,\mathsf{t}_{M-1}$ are as in (4.14). Hence,
@@
where we are using the fact that $\mathsf{t}_{k}\geq 0$ and $y_{\mathsf{t}_{0}}\geq y_{\mathsf{t}_{k}}-\delta+1$ for all $y_{\mathsf{t}_{k}}-y_{\mathsf{t}_{0}}+1\leq\delta$. Thus, we conclude that the right-hand side of (4.23) is bounded above by
@@
By using the bound from (4.28) in (4.26) and summing over $k$, we arrive at the desired estimate (4.18).
@@
(a) Now we prove part (a). We start with rewriting $\chi_{\mathfrak{h}}(\eta,\xi)$ under the assumption that $\mathfrak{h}\in m\mathrm{NW}_{0}$, recall the definition of $\chi_{\mathfrak{h}}(\eta,\xi)$ from (2.9). For $\mathrm{supp}(\mathfrak{h})=\{\omega_{0},\ldots,\omega_{M-1}\}$ with $0=\omega_{0}>\cdots>\omega_{M-1}$, one has $\mathbb{P}(\bm{\tau}_{+}\neq 0)=0$ where $\bm{\tau}_{+}$ is defined in (2.16). Thus, it is easy to check that the first and third term on the right-hand side of (2.9) are both equal to $\frac{\mathrm{e}^{\mathfrak{h}(0)(\eta-\xi)}}{\eta-\xi}$ and they cancel each other. On the other hand, since $\mathbb{P}(\bm{\tau}_{-}\notin\{\omega_{0},\ldots,\omega_{M-1}\})=0$, the second term on the right-hand side of (2.9) is given by
@@
In order to rigorously show the above convergence, we need to show that (4.38) is uniformly bounded and the dominated convergence theorem applies. Note that the right-hand side of (4.38) is the same as that of (4.21). Therefore, (4.19) gives a uniform bound for (4.38). This completes the proof. ∎
@@ 4.3 Convergence of the series expansion
In this section we prove Lemma 4.3 and 4.4. We will make the additional assumption that $0<\tau_{1}<\cdots<\tau_{m}$ to make the presentation lighter. The convergence results and arguments in this section still work if $\tau_{i}=\tau_{i+1}$ for some $i$ but the contours need to be chosen carefully to make sure that the integrand has the desired super-exponential decay. Alternatively, one can directly work with the limiting KPZ fixed point formula which is continuous with respect to the limit $\tau_{i+1}\to\tau_{i}$ and our choices of the angles in Figure 1 guarantees the convergence of the formula (2.21) even when some time parameters are equal.
@@
Thus, for such $\eta$ we have
@@
For the first part of the claim, note that the integrand on the right-hand side of (4.50) converges pointwise to the integrand on the right-hand side of (2.21) by Lemma 4.6(a) and Proposition 4.2(a). On the other hand, the cubic exponential decay bound for $|\mathrm{f}_{i}^{\varepsilon}|$ from Lemma 4.6(b) and the quadratic exponential growth estimate (4.19) implies that the integrand on the right-hand side of (4.50) has a cubic exponential decay in every variable $\eta_{i_{\ell}}^{(\ell)},\xi_{i_{\ell}}^{(\ell)}$ as they go to $\infty$, uniform in $\varepsilon$. Thus, the dominated convergence theorem applies, and the first claim is proved.
For the second part of the claim, note that on $A^{c}$, at least one of the variables, say $\xi_{1}^{(1)}$, lies outside $\widehat{\Omega}_{0}^{\varepsilon}$. Then Lemma 4.6(b) implies that $|\mathrm{f}_{1}^{\varepsilon}(\xi_{1}^{(1)})|\leq C\exp(-c\varepsilon^{-\frac{3}{2}})$ for all $\xi_{1}^{(1)}\in{\Gamma}_{1,\mathrm{L}}^{\varepsilon}\backslash\widehat{\Omega}_{0}^{\varepsilon}$. On the other hand, Proposition 4.2(b) implies that
@@
for some $c,C>0$ and all $1\leq j\leq n_{1}$. The other parts of the integrand remain bounded. Thus,
@@
This proves the second part of the claim, and Lemma 4.3 follows.
@@
for some constant $c>0$ independent of $\varepsilon$ and $\bm{\xi}^{(1)},\bm{\eta}^{(1)}$. Thus, by Hadamard’s inequality
@@
for $1\leq\ell\leq m-1$. Here $D$ is chosen to be the reciprocal of the minimal distance between different $\Gamma^{\varepsilon}$ contours, which can be made positive. Thus,
@@ 5 From multiple narrow wedges to compactly supported initial conditions
In this section we extend the multipoint formula (1.6) to any $\mathfrak{h}\in\mathrm{UC}_{c}$ using a density argument. We choose to work at the level of the KPZ fixed point formula, which enjoys more symmetry and nicer decay properties.
@@ 5.1 Density and Approximation
Our starting point is the following proposition, asserting that our formula is continuous with respect to the initial condition on the space $\mathrm{UC}_{c}$.
@@
for all $\xi\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$. Recall the kernel $\chi_{\mathfrak{h}}$ was defined in Definition 2.2, equation (2.9). Consequently,
$\lim_{n\to\infty}\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}^{n}}(z_{1},\ldots,z_{m-1})=\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}).$ (5.2)
@@
This completes the proof of (5.1). (5.2) follows from the dominated convergence theorem and a similar bound as in Lemma 4.4 whose proof is almost identical, and we omit it here. ∎
To extend (1.6) to all $\mathfrak{h}\in\mathrm{UC}_{c}$ using Proposition 5.1, we also need the following proposition, asserting that the space $m\mathrm{NW}$ is dense in $\mathrm{UC}_{c}$.
@@
It suffices to consider the restrictions of the functions on [L,L][-L,L]. For each $n$ consider the dyadic intervals $I_{n,k}:=[\frac{k}{2^{n}}\cdot L,\frac{k+1}{2^{n}}\cdot L]$, for $k=-2^{n},-2^{n}+1,\ldots,2^{n}-1$. On each interval $I_{n,k}$ the maximum of $\mathfrak{h}$ exists, let $m_{n,k}\in I_{n,k}$ be one of the argmax, and set
@@
Therefore, $\mathfrak{h}^{n}\to\mathfrak{h}$ locally in $\mathrm{UC}$, and hence globally since $\mathfrak{h}^{n},\mathfrak{h}$ are supported in [L,L][-L,L]. ∎
@@
Recall that what we have shown in Proposition 5.2 is that the formula (1.6) holds for initial conditions $\mathfrak{h}\in m\mathrm{NW}_{0}$. We would like to prove it for $\mathfrak{h}\in m\mathrm{NW}$ by a shift argument. To this end it is more convenient to have a more general version of the characteristic function $\chi_{\mathfrak{h}}$, defined through (2.15), instead of the original (2.9). The goal of this section is to prove the equivalence of the two, i.e., Proposition 2.4. To this end, we denote
@@
$\quad+\mathrm{e}^{\omega\eta^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{+s\eta}\cdot\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(-\bm{\tau}^{\omega}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}^{\omega}_{+})\xi\right)\mathbf{1}_{\bm{\tau}^{\omega}_{+}<\infty}\right]$
$\quad-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(\bm{\tau}_{-}^{\omega}\eta^{2}+\mathbf{B}(\bm{\tau}^{\omega}_{-})\eta-\bm{\tau}_{+}^{\omega}\xi^{2}-\mathbf{B}(\bm{\tau}^{\omega}_{+})\xi\right)\mathbf{1}_{|\bm{\tau}^{\omega}_{\pm}|<\infty}\right]$
@@
where $\omega_{1}>\cdots>\omega_{k}$. To lighten the notations, we will fix $\eta,\xi$ and denote
@@
$\chi_{\mathfrak{h}}^{\omega,1}(\eta,\xi)$ $=\mathrm{e}^{-\omega\xi^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\int_{\mathbb{R}}\mathop{}\!\mathrm{d}b_{1}\,\mathrm{e}^{-s\xi}\mathrm{p}_{\omega-\omega_{1}}(b_{1}-s)\,E^{-}(\omega_{1},b_{1})$
$=\mathrm{e}^{-\omega_{1}\xi^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}b_{1}\,\mathrm{e}^{-b_{1}\xi}\,E^{-}(\omega_{1},b_{1})=\chi_{\mathfrak{h}}^{\omega_{1},1}(\eta,\xi),$
@@
Recall that $\mathrm{p}_{t-s}(x-y)$ is the transition density of a Brownian motion with diffusivity constant $2$. On the other hand,
@@
$\quad+\mathbf{1}_{\omega=\omega_{i+1}}\frac{\mathrm{e}^{\omega_{i+1}(\eta^{2}-\xi^{2})+\theta_{i+1}(\eta-\xi)}}{\eta-\xi}.$
By combining (5.5), (5.6) and (5.7), we conclude that
@@
for all $\omega\in[\omega_{i+1},\omega_{i})$. Thus, by induction $\chi_{\mathfrak{h}}^{\omega}=\chi_{\mathfrak{h}}$ for all $\omega\in\mathbb{R}$ and $\mathfrak{h}\in m\mathrm{NW}$.
For general $\mathfrak{h}\in\mathrm{UC}_{c}$ supported on [L,L][-L,L], bounded above by $\beta$, we use Proposition (5.2) to find a sequence $\{\mathfrak{h}^{n}\}_{\geq 1}\subset m\mathrm{NW}$ bounded above by $\beta$ whose supports are contained in [L,L][-L,L] , such that $\mathfrak{h}^{n}\to\mathfrak{h}$ in $\mathrm{UC}$. Then by Proposition 5.1 we know $\chi_{\mathfrak{h}^{n}}\to\chi_{\mathfrak{h}}$ as $n\to\infty$. A minor variant of Proposition 5.1 with the same proof shows that $\chi^{\omega}_{\mathfrak{h}^{n}}\to\chi^{\omega}_{\mathfrak{h}}$ as $n\to\infty$ for any $\omega\in\mathbb{R}$ as well. Thus,
@@
Recall that in Proposition 4.2, we have shown Theorem 1.2 for $\mathfrak{h}\in m\mathrm{NW}_{0}$. We first extend it to $\mathfrak{h}\in m\mathrm{NW}$ using Proposition 2.4. There exists $\alpha,\beta\in\mathbb{R}$ such that $\mathfrak{h}^{\alpha,\beta}:=\mathfrak{h}(\cdot+\alpha)+\beta\in m\mathrm{NW}_{0}$. Thus, by the invariance property of the KPZ fixed point (see, e.g., [MQR21][Theorem 4.5]) we have
@@
We apply Proposition 4.2 to the right-hand side of (5.8) and compare the resulting formula with the right-hand side of (1.6). In order to prove Theorem 1.2 holds for $\mathfrak{h}\in m\mathrm{NW}$, it suffices to show that